OpenStudy (anonymous):

Can someone please help me with this? Neveah has created the function f(x) = the quantity of 2x plus 4, divided by 3 to represent the growth of her hair, where x represents the number of months since her last hair cut. Neveah discovers that using the inverse function to solve for x = 6, she can predict when her hair will have grown 6 inches. Explain to Neveah how to accomplish this, using complete sentences.

OpenStudy (ranga):

Is f(x) = (2x + 4) / 3 or 2x + 4/3 ?

OpenStudy (anonymous):

f(x) = (2x + 4) / 3

OpenStudy (ranga):

f(x) = (2x + 4) / 3 Find the inverse function. Step 1: Rewrite the equation, by using y in the place of f(x): y = (2x + 4) / 3 Step 2: To find the inverse function, switch x and y. That is put x in the place of y and y in the place of x: x = (2y + 4) / 3 Step 3: Solve for y: x = (2y + 4) / 3 multiply both sides by 3 3x = 2y + 4 subtract 4 3x - 4 = 2y divide by 2 (3x - 4) / 2 = y The inverse function is: f^-1(x) = (3x - 4) / 2

OpenStudy (ranga):

To find out how many months it will take for her hair to grow 6 inches, put x = 6 in the inverse function: ( 3 * 6 - 4 ) / 2 = (18 - 4) / 2 = 14 / 2 = 7 months.

OpenStudy (anonymous):

Okay, thank you so much for clearing this up for me. You've been a big help.

OpenStudy (ranga):

You are welcome.